OpenStudy (anonymous):

y/z divide z/t =1 for y solve each formula for the given letter

OpenStudy (anonymous):

my computer have a problem don't let me post

OpenStudy (anonymous):

\[(\frac{y}{z})\div(\frac{z}{t})=1\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

cross and multiply u get \[\frac{yt}{z ^{2}}=1\]

OpenStudy (anonymous):

\[yt=1\times z ^{2}\]\[y=\frac{z ^{2}}{t}\]